Earl Ezekiel wants to retire in San Diego when he is 65 years old. Earl is now 55. He believes he will need $410,000 to retire comfortably. To date, Earl has set aside no retirement money. Assume Earl gets 4% interest compounded semiannually.
  
How much must Earl invest today to meet his $410,000 goal?(Do not round intermediate calculations. Round your answer to the nearest cent.)
Answer #1

Present value = Future value / (1+( rate per annum/compounding per year))^ (no. of year*compounding per year)

= 410000 /(1+(.04/2))^(10*2)

= 410000/1.02^20

= 410000/1.48594739598

= $275918.25